OpenStudy (anonymous):

Simplify: w^14/w^7*w^15/w^5 w^6 w^70 w^5 w^17

OpenStudy (anonymous):

anybody

OpenStudy (anonymous):

@zpupster

OpenStudy (anonymous):

reframe your question: if it is \[\frac{ W ^{14} }{W^{7}} * \frac{ W^{15} }{ W^{5} }\] then your answer is: \[= W ^{7} \times W ^{10} = W ^{7 + 10} = W ^{17}\] Note:Since,the base are same therefore the multiplication results in the additions of the powers